Use boolean algebra to reduce; qm+q!p+q!m+!q!m+!mp
Answer #1
qm+q!p+q!m+!q!m+!mp
=qm+q!m+q!p+q!m+!q!m+!mp
= qm+q!m+q!p+(q+!q)!m+!mp
= qm+q!m+q!p+(1)!m+!mp
= qm+q!m+q!p+!m+!mp
= qm+q!m+q!p+!m(1+p)
= qm+q!m+q!p+!m(1)
= qm+q!m+q!p+!m
= q(m+!m)+q!p+!m
= q(1)+q!p+!m
= q+q!p+!m
= q(1+!p)+!m
= q(1)+!m
= q+!m